Crew Chief Shift Log, 2025-05-05 DAY
Start of shift: Beam to B
Start of shift: Beam to B
After Yan made some injector changes. He wanted to check how the halo rates in Hall B would respond. Spoke with the Hall as we told them beam would be off for a longer period of time than they were expecting (sorry!). The Hall agreed to accept some beam to compare halo rates. Starting off at 5 nA and ramping up to 325 nA.
For the upcoming maintenance day we would like to repair some damaged power and signal cables in the North Linac for the BLM's. This would require access to the NL if possible.
We will also install the HV supply to FSD system interface box that we removed earlier for repairs (reference log #4356378.) and has been running on our test stand for 3 weeks with know errors.